OpenStudy (anonymous):

I can't remember the steps to do this problem: (-4x)^2(3x^8)^2

OpenStudy (anonymous):

PEMDAS is needed here. So, first, square the terms in the parenthesis.

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

-16x and 9x^8?

OpenStudy (anonymous):

square the entire term (-4x) and you get....

OpenStudy (saifoo.khan):

\[\Large (-4x)(-4x) \times (3x^8)(3x^8)\] Can u solve now?

OpenStudy (anonymous):

Ah! Lemme give it a shot!

OpenStudy (anonymous):

16x * 9x^16?

OpenStudy (anonymous):

Almost. What is x*x?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

Same rule as when you did (x^8) * (x^8)

OpenStudy (anonymous):

So the answer is 144x^8 ?

OpenStudy (anonymous):

Sorry, 144x^16

OpenStudy (saifoo.khan):

144x^18

OpenStudy (anonymous):

How do you get 18 as an exponent?

OpenStudy (saifoo.khan):

I am sure u missed to add the "left exponents"

OpenStudy (saifoo.khan):

(-4x)(-4x) -4 * - 4 = 16 x * x = x^2

OpenStudy (anonymous):

Oh! Yes I did miss that part! I see now and thank you all so very much :)

OpenStudy (saifoo.khan):

Welcome.